Can you help me with this question, please

9. A Normal body temperature is 98.60F. (Show your work)

a.Calculate the corresponding temperature in degree Celsius and then in Kelvin units.

b. When you enter a swimming pool, the water may feel quite cold. After a while, though, your body "gets used to it," and the water no longer feels so cold. Use the concept of heat to explain what is going on. (note that heat flows because of a temperature difference.)
Answer #1

a) T=98.6 C = 5x (F-32) - 12 - . . C = 5x (98.6-32) .37°C K = c+273 K = 37+2 73 - 310 Temperature is 37°C OR 31ok (6) When weour body and water and the watel no longer feels so Cold